OpenStudy (anonymous):

Hardest math day!!! Hope you can answer it :D One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

jimthompson5910 (jim_thompson5910):

Let x = # of humans and y = # of horses There are 74 heads. Since both have one head each, we know x+y = 74 Humans have 2 legs while horses have 4, so 2x+4y = 196 To find the answer, simply solve this system.

OpenStudy (anonymous):

i do know the answer...i just want everyone to have answer and have fun :D

OpenStudy (anonymous):

are there any siamese horses?

OpenStudy (across):

This is really fun.

OpenStudy (anonymous):

haha maybe yes :D

OpenStudy (anonymous):

yes i agree with @across

OpenStudy (anonymous):

well here is the answer :D Let's assume that HM = Human and HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50 So, the solution is 24 horses and 50 humans. :D